# A Free Probabilistic Framework for Denoising Diffusion Models: Entropy, Transport, and Reverse Processes

### v1

@@ VII-C On the $\Gamma$-limit of the free JKO functionals
which establishes the lower semicontinuity in item 1. The argument is the free-probability analogue of the classical lower semicontinuity of $W_{2}$ on $\mathcal{P}_{2}(\mathbb{R})$, see also [31, Ch. 2].
@@
where $\lambda_{A}$ denotes the empirical spectral law of $A$. Let $\nu_{n}\rightharpoonup\nu$ narrowly with $\sup_{n}\int x^{2}\,d\nu_{n}(x)<\infty$. For each fixed $R>0$, the sets of $m\times m$ microstates $\Gamma_{m}(\nu_{n};R,\varepsilon)$ corresponding to $\nu_{n}$ converge in the Hausdorff sense (with respect to the normalized trace topology) to $\Gamma_{m}(\nu;R,\varepsilon)$ as $n\to\infty$, because narrow convergence with uniformly bounded second moments implies convergence of all *-moments of bounded degree. By the monotonicity of $\mathrm{Vol}[\Gamma_{m}(\nu;R,\varepsilon)]$ in $R$ and $\varepsilon$, and the subadditivity of $\frac{1}{m^{2}}\log\mathrm{Vol}$ in $m$, one can apply a diagonal argument in $(m,R,\varepsilon)$ to pass to the limit (see [34, Thm. 6.5.1] and [33]), obtaining
@@
(iii) Consequences..  Since the free energy functional $\mathcal{F}[\nu]\mathrel{\mathop{\ordinarycolon}}=\tfrac{1}{2}W_{2,\mathrm{free}}^{2}(\nu,\mu_{\mathrm{sc}})-\chi[\nu]$ is the sum of a lower semicontinuous and an upper semicontinuous term, it is itself lower semicontinuous on $\mathcal{P}_{2}^{\mathrm{free}}$. The moment bound guarantees tightness of sublevel sets of $\mathcal{F}$, which are therefore relatively compact in the narrow topology. This completes the proof. ∎
@@
We will prove the three items in order. The proof strategy is standard (cf. Ambrosio–Gigli–Savaré [31, Ch. 4]), but we carefully check each step in the free setting using Lemma 4.
@@
Theorems 2 ($\Gamma$-convergence of $\mathcal{F}^{\tau}$) and standard minimizing-movement compactness arguments (see Ambrosio–Gigli–Savaré [31, Ch. 4]) imply the following.
@@
Let $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$ denote the space of compactly supported spectral measures with finite second moment, endowed with the free $2$-Wasserstein distance of Biane–Voiculescu. Assume that the free energy functional
@@
If $\mathcal{F}$ is $\lambda$-geodesically convex in $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$, then standard variational arguments (see [31, Ch. 4]) yield the differential contractivity estimate
@@
This holds, for instance, for the quadratic potential plus free entropy under compact support (see the discussion in [34, Sec. 6] and [16].
@@
In the present free setting, using the continuity equation representation and Lemma 4, one identifies (cf. [31, Ch. 10])
@@
The standard minimizing-movement compactness and a priori bounds from [31, Chs. 2–3 & 4] apply verbatim by Lemma 4, yielding (up to subsequences) $\mu^{\tau}(t)\rightharpoonup\mu_{t}$ and the discrete energy dissipation estimate
@@
Summing over $k$ and passing to the limit gives the EDI (25). Under (VII), the AGS theory yields the EVI (26) and uniqueness of the gradient flow [31, Thm. 4.0.4 & Ch. 11]. The identification (ID) follows from the metric characterization of slopes via the continuity equation and the free integration-by-parts identity defining $\Xi_{t}$; hence, the velocity field is $b_{t}=-\frac{1}{2}x-\Xi_{t}$, which gives (27). By Biane–Speicher free Itô theory [14], (27) is equivalent to the forward free SDE (28); its time reversal is exactly Theorem 1, where the additional $\Xi_{t}$ term appears as the free score correction in the backward drift. ∎
@@
If $t\mapsto\mu_{t}$ is a curve of maximal slope for $\mathcal{F}$ (i.e., equality holds in (25)), then the metric derivative and slope coincide a.e.: $|\mu^{\prime}|(t)=|\partial\mathcal{F}|(\mu_{t})=\|-\tfrac{1}{2}x-\Xi_{t}\|_{L^{2}(\mu_{t})}$, which is precisely the Fisher-information dissipation identity in the free setting; see [31, Ch. 10] and [34, Sec. 6].
@@ VIII Information-Theoretic Interpretation
This section provides the detailed derivations of the free de Bruijn identity, the Fisher–information flow, and the associated HWI-type inequalities that underlie the information-theoretic interpretation developed in Section VI. The arguments presented here are technical and are included for completeness. They also form the analytical basis for the summary given in Appendix D. The forward free diffusion process defined in (7) admits a natural interpretation in terms of non-commutative information measures. In classical information theory, Gaussian heat flow increases Shannon entropy and satisfies the de Bruijn identity, which relates the derivative of the entropy to the Fisher information. The same principle holds in the free setting, with Voiculescu’s free entropy $\chi$ playing the role of Shannon entropy and the free Fisher information $\Phi^{*}$ taking the role of classical Fisher information.
@@ VIII-A Free de Bruijn Identity
and $\Xi_{s}$ is the conjugate variable of $X_{s}$. Equation (29) is the non-commutative analogue of the classical de Bruijn identity and shows that the forward free diffusion is an entropy-increasing flow. In particular, $\chi(X_{s})$ is non-decreasing in $s$, with strict monotonicity unless $X_{s}$ is already semicircular. Equation (30) shows that the forward free diffusion monotonically increases Voiculescu’s free entropy, driving the spectral law toward the semicircular equilibrium. This evolution is illustrated in Fig. 1, where the forward process (top) broadens structured spectra while the reverse transport (bottom) performs free deconvolution, recovering low-entropy spectral states.
@@ VIII-B Free Score Function and Reverse Drift
The conjugate variable $\Xi_{s}$ plays the same role in free probability as the score function $\nabla\log p(x)$ in classical statistics. It represents the $L^{2}$-gradient of the log non-commutative density and encodes how the spectral law of $X_{s}$ deviates from the equilibrium semicircular law. The reverse-time dynamics of Theorem 1 can thus be viewed as a free analogue of the score-based generative modeling framework.
@@
For a single self-adjoint non-commutative random variable, Voiculescu’s free entropy has the variational derivative
@@ VIII-C Entropy Production and Thermodynamic Inequalities
and the free Fisher information is connected to the derivative of $\chi$ through the non-commutative analogue of the classical de Bruijn identity. The free analog of the Fisher information inequality, due to Biane and Voiculescu, states that for free convolution with a semicircular element $S_{t}$,
@@ IX-E Consequences
Thus, the free diffusion model satisfies a hierarchy of functional inequalities entirely analogous to the classical Gaussian case, now formulated in non-commutative Wasserstein geometry. This establishes the complete correspondence between free stochastic dynamics, functional inequalities, and information geometry.
@@ X-B Approximation and Universality of the Reverse Process
defines a generative process. The classical theory of Schrödinger bridges, together with its non-commutative analogue, shows that for any absolutely continuous target law $\mu_{0}$, there exists a control $U_{t}=-\Xi_{t}$ that transports $\mu_{T}$ back to $\mu_{0}$.
@@ XI Conclusion
$W_{2,\mathrm{free}}$ Free 2-Wasserstein distance on $\mathcal{P}_{2}^{\mathrm{free}}$
@@ Appendix A PDE-based derivation of Theorem 1
The reverse-time SDE can also be obtained from the PDE governing the evolution of the spectral law of $X_{t}$ under the forward SDE. This derivation requires no stochastic calculus beyond identifying the conjugate variable with the non-commutative gradient of the log-density.
Let $\mu_{t}\mathrel{\mathop{\ordinarycolon}}=\mu_{X_{t}}$ denote the spectral measure of $X_{t}$ for $t\in[0,T]$. Under the forward SDE (LABEL:(F)), $\mu_{t}$ evolves according to the free Fokker–Planck equation
@@ Appendix B Worked example: Dirac initial law (explicit forward and reverse flows)
This closed-form example thus verifies the abstract formulas of Theorem 1: the conjugate variable appears as the non-commutative score, and it yields an explicit analytic expression for the reverse-time drift in the freely Gaussian (semicircular) case. ∎
@@ Appendix C Derivation of the Free Clark–Ocone Formula and Time–Reversal Identity
This appendix presents the analytic foundation underlying Theorem 1. We first derive the free Clark–Ocone decomposition—the noncommutative analogue of the classical martingale representation theorem—and then justify the time–reversal identity used to obtain the reverse–time free SDE. Throughout, $(\mathcal{A},\varphi)$ denotes a tracial $W^{\ast}$–probability space, and $(S_{t})_{t\geq 0}$ a free Brownian motion (semicircular process) with forward filtration $\mathcal{F}_{t}=\sigma\{S_{u}\mathrel{\mathop{\ordinarycolon}}0\leq u\leq t\}$.
@@ C.1  Functional–Analytic Framework
C.2  Free Clark–Ocone Decomposition
Let $\mathcal{H}_{\mathrm{pred}}$ denote the closed subspace of $L^{2}([0,T]\times\mathcal{A}\!\otimes\!\mathcal{A})$ consisting of predictable (i.e. $\mathcal{F}_{t}$–measurable) processes. For $F\in\mathrm{Dom}(D)$, define the predictable projection
$u_{t}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{\varphi}[D_{t}F\mid\mathcal{F}_{t}],$
where $\mathbb{E}_{\varphi}[\cdot\mid\mathcal{F}_{t}]$ denotes the orthogonal projection of $L^{2}(\mathcal{A},\varphi)$ onto $L^{2}(\mathcal{F}_{t},\varphi)$. Then $u=(u_{t})_{t}$ is the unique element of $\mathcal{H}_{\mathrm{pred}}$ minimizing $\|D_{t}F-v_{t}\|_{L^{2}}$ over all predictable $v_{t}$; that is, $u$ is the orthogonal projection of $DF$ onto $\mathcal{H}_{\mathrm{pred}}$.
Representation theorem.
Combining (52) with the orthogonality of the projection, for all $G\in\mathrm{Dom}(\delta)$,
$\varphi[F\,\delta(G)]=\int_{0}^{T}\varphi[\langle D_{t}F,\,G_{t}\rangle_{\mathrm{free}}]\,\mathrm{d}t=\int_{0}^{T}\varphi[\langle u_{t},\,G_{t}\rangle_{\mathrm{free}}]\,\mathrm{d}t=\varphi\!\left[\Big(\!\int_{0}^{T}u_{t}\,\#\,\mathrm{d}S_{t}\!\Big)G\right].$
Since $\mathrm{Dom}(\delta)$ is dense in $L^{2}(\mathcal{A},\varphi)$, the representing element is unique, and we obtain the free Clark–Ocone formula
$F=\mathbb{E}_{\varphi}[F]+\int_{0}^{T}\mathbb{E}_{\varphi}[D_{t}F\mid\mathcal{F}_{t}]\,\#\,\mathrm{d}S_{t},\qquad F\in W^{1,2}_{\mathrm{free}}.$ (53)
@@ C.2  Free Clark–Ocone Decomposition
  3. (iii)
    Uniqueness of the integrand follows from the isometry (51): if two predictable processes $u,v$ satisfy (53), then $\int_{0}^{T}(u_{t}-v_{t})\,\#\,\mathrm{d}S_{t}=0$ implies $u=v$ in $L^{2}$.
@@ C.3  Free Time–Reversal of Semicircular Diffusions
The rigorous time–reversal step in Theorem 1 follows from Biane’s duality theorem for free stochastic integrals [29, Proposition 3.1]. For any adapted process $U_{t}\in L^{2}(\mathcal{A},\varphi)$ define
@@ D-A Voiculescu’s Free Entropy Functional
For a self-adjoint non-commutative random variable $X$ with spectral distribution $\mu$, Voiculescu’s free entropy $\chi(\mu)$ is defined by
@@ D-D Euler–Lagrange Equation and Free McKean–Vlasov Flow
The free entropy $\chi(\mu)$ can be regarded as the non-commutative analogue of the classical McKean–Vlasov free energy functional

### v2

@@ VII-C On the $\Gamma$-limit of the free JKO functionals
which establishes the lower semicontinuity in item 1. The argument is the free-probability analogue of the classical lower semicontinuity of $W_{2}$ on $\mathcal{P}_{2}(\mathbb{R})$, see also [38, Ch. 2].
@@
where $\lambda_{A}$ denotes the empirical spectral law of $A$. Let $\nu_{n}\rightharpoonup\nu$ narrowly with $\sup_{n}\int x^{2}\,d\nu_{n}(x)<\infty$. For each fixed $R>0$, the sets of $m\times m$ microstates $\Gamma_{m}(\nu_{n};R,\varepsilon)$ corresponding to $\nu_{n}$ converge in the Hausdorff sense (with respect to the normalized trace topology) to $\Gamma_{m}(\nu;R,\varepsilon)$ as $n\to\infty$, because narrow convergence with uniformly bounded second moments implies convergence of all *-moments of bounded degree. By the monotonicity of $\mathrm{Vol}[\Gamma_{m}(\nu;R,\varepsilon)]$ in $R$ and $\varepsilon$, and the subadditivity of $\frac{1}{m^{2}}\log\mathrm{Vol}$ in $m$, one can apply a diagonal argument in $(m,R,\varepsilon)$ to pass to the limit (see [41, Thm. 6.5.1] and [40]), obtaining
@@
(iii) Consequences.  Since the free energy functional $\mathcal{F}[\nu]\mathrel{\mathop{\ordinarycolon}}=\tfrac{1}{2}W_{2,\mathrm{free}}^{2}(\nu,\mu_{\mathrm{sc}})-\chi[\nu]$ is the sum of a lower semicontinuous and an upper semicontinuous term, it is itself lower semicontinuous on $\mathcal{P}_{2}^{\mathrm{free}}$. The moment bound guarantees tightness of sublevel sets of $\mathcal{F}$, which are therefore relatively compact in the narrow topology. This completes the proof. ∎
@@
We will prove the three items in order. The proof strategy is standard (cf. Ambrosio–Gigli–Savaré [38, Ch. 4]), but we carefully check each step in the free setting using Lemma 4.
@@
Theorems 2 ($\Gamma$-convergence of $\mathcal{F}^{\tau}$) and standard minimizing-movement compactness arguments (see Ambrosio–Gigli–Savaré [38, Ch. 4]) imply the following.
@@
Let $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$ denote the space of compactly supported spectral measures with finite second moment, endowed with the free $L_{2}$-Wasserstein distance of Biane–Voiculescu. Assume that the free energy functional
@@
If $\mathcal{F}$ is $\lambda$-geodesically convex in $(\mathcal{P}_{2}^{\mathrm{free}},W_{2,\mathrm{free}})$, then standard variational arguments (see [38, Ch. 4]) yield the differential contractivity estimate
@@
This holds, for instance, for the quadratic potential plus free entropy under compact support (see the discussion in [41, Sec. 6] and [16].
@@
In the present free setting, using the continuity equation representation and Lemma 4, one identifies (cf. [38, Ch. 10])
@@
The standard minimizing-movement compactness and a priori bounds from [38, Chs. 2–3 & 4] apply verbatim by Lemma 4, yielding (up to subsequences) $\mu^{\tau}(t)\rightharpoonup\mu_{t}$ and the discrete energy dissipation estimate
@@
Summing over $k$ and passing to the limit gives the EDI (25). Under (VII), the AGS theory yields the EVI (26) and uniqueness of the gradient flow [38, Thm. 4.0.4 & Ch. 11]. The identification (ID) follows from the metric characterization of slopes via the continuity equation and the free integration-by-parts identity defining $\Xi_{t}$; hence, the velocity field is $b_{t}=-\frac{1}{2}x-\Xi_{t}$, which gives (27). By Biane–Speicher free Itô theory [14], (27) is equivalent to the forward free SDE (28); its time reversal is exactly Theorem 1, where the additional $\Xi_{t}$ term appears as the free score correction in the backward drift. ∎
@@
If $t\mapsto\mu_{t}$ is a curve of maximal slope for $\mathcal{F}$ (i.e., equality holds in (25)), then the metric derivative and slope coincide a.e.: $|\mu^{\prime}|(t)=|\partial\mathcal{F}|(\mu_{t})=\|-\tfrac{1}{2}x-\Xi_{t}\|_{L^{2}(\mu_{t})}$, which is precisely the Fisher-information dissipation identity in the free setting; see [38, Ch. 10] and [41, Sec. 6].
@@ VIII Information-Theoretic Interpretation
This section provides the detailed derivations of the free de Bruijn identity, the Fisher–information flow, and the associated HWI-type inequalities that underlie the information-theoretic interpretation developed in Section VI. The arguments presented here are technical and are included for completeness. They also form the analytical basis for the summary given in Appendix D. The forward free diffusion process defined in (7) admits a natural interpretation in terms of noncommutative information measures. In classical information theory, Gaussian heat flow increases Shannon entropy and satisfies the de Bruijn identity, which relates the derivative of the entropy to the Fisher information. The same principle holds in the free setting, with Voiculescu’s free entropy $\chi$ playing the role of Shannon entropy and the free Fisher information $\Phi^{*}$ taking the role of classical Fisher information.
@@ VIII-A Free de Bruijn Identity
and $\Xi_{s}$ is the conjugate variable of $X_{s}$. Equation (29) is the noncommutative analogue of the classical de Bruijn identity and shows that the forward free diffusion is an entropy-increasing flow. In particular, $\chi(X_{s})$ is non-decreasing in $s$, with strict monotonicity unless $X_{s}$ is already semicircular. Equation (30) shows that the forward free diffusion monotonically increases Voiculescu’s free entropy, driving the spectral law toward the semicircular equilibrium. This evolution is illustrated in Fig. 1, where the forward process (top) broadens structured spectra while the reverse transport (bottom) performs free deconvolution, recovering low-entropy spectral states.
@@ VIII-B Free Score Function and Reverse Drift
The conjugate variable $\Xi_{s}$ plays the same role in free probability as the score function $\nabla\log p(x)$ in classical statistics. It represents the $L^{2}$-gradient of the log noncommutative density and encodes how the spectral law of $X_{s}$ deviates from the equilibrium semicircular law. The reverse-time dynamics of Theorem 1 can thus be viewed as a free analogue of the score-based generative modeling framework.
@@
For a single self-adjoint noncommutative random variable, Voiculescu’s free entropy has the variational derivative
@@ VIII-C Entropy Production and Thermodynamic Inequalities
and the free Fisher information is connected to the derivative of $\chi$ through the noncommutative analogue of the classical de Bruijn identity. The free analog of the Fisher information inequality, due to Biane and Voiculescu, states that for free convolution with a semicircular element $S_{t}$,
@@ IX-E Consequences
Thus, the free diffusion model satisfies a hierarchy of functional inequalities entirely analogous to the classical Gaussian case, now formulated in noncommutative Wasserstein geometry. This establishes the complete correspondence between free stochastic dynamics, functional inequalities, and information geometry.
@@ X-B Approximation and Universality of the Reverse Process
defines a generative process. The classical theory of Schrödinger bridges, together with its noncommutative analogue, shows that for any absolutely continuous target law $\mu_{0}$, there exists a control $U_{t}=-\Xi_{t}$ that transports $\mu_{T}$ back to $\mu_{0}$.
@@ XI Conclusion
$W_{2,\mathrm{free}}$ Free $L_{2}$-Wasserstein distance on $\mathcal{P}_{2}^{\mathrm{free}}$
@@ Appendix A PDE-based derivation of Theorem 1
The reverse-time SDE can also be obtained from the PDE governing the evolution of the spectral law of $X_{t}$ under the forward SDE. This derivation requires no stochastic calculus beyond identifying the conjugate variable with the noncommutative gradient of the log-density.
Let $\mu_{t}\mathrel{\mathop{\ordinarycolon}}=\mu_{X_{t}}$ denote the spectral measure of $X_{t}$ for $t\in[0,T]$. Under the forward SDE (7), $\mu_{t}$ evolves according to the free Fokker–Planck equation
@@ Appendix B Worked example: Dirac initial law (explicit forward and reverse flows)
This closed-form example thus verifies the abstract formulas of Theorem 1: the conjugate variable appears as the noncommutative score, and it yields an explicit analytic expression for the reverse-time drift in the freely Gaussian (semicircular) case. ∎
@@ Appendix C Derivation of the Free Clark–Ocone Formula and Time–Reversal Identity
This appendix presents the analytic foundation underlying Theorem 1. We first derive and prove Lemma 1—the free Clark–Ocone decomposition, a noncommutative analogue of the classical martingale representation theorem—and then justify the time–reversal identity used to obtain the reverse–time free SDE. Throughout, $(\mathcal{A},\varphi)$ denotes a tracial $W^{\ast}$–probability space, and $(S_{t})_{t\geq 0}$ a free Brownian motion (semicircular process) with forward filtration $\mathcal{F}_{t}=\sigma\{S_{u}\mathrel{\mathop{\ordinarycolon}}0\leq u\leq t\}$.
@@ C.1  Functional–Analytic Framework
C.2  Proof of Lemma 1 (Free Clark–Ocone Formula)
The operator $D$ is densely defined and closable on $L^{2}(\mathcal{A},\varphi)$, and $\delta=D^{\ast}$ by (52). For $F\in\mathrm{Dom}(D)$, define the predictable projection $u_{t}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{\varphi}[D_{t}F\mid\mathcal{F}_{t}]$. Then $u\in L^{2}_{\mathrm{ad}}([0,T];\mathcal{A})$ and $\delta(u)=\int_{0}^{T}u_{t}\,\#\,\mathrm{d}S_{t}$ is well defined. Using orthogonality of the conditional expectation and the adjointness relation, one obtains
$\varphi[F\,\delta(G)]=\int_{0}^{T}\varphi[\langle D_{t}F,G_{t}\rangle_{\mathrm{free}}]\,\mathrm{d}t=\int_{0}^{T}\varphi[\langle u_{t},G_{t}\rangle_{\mathrm{free}}]\,\mathrm{d}t=\varphi\!\left[\!\Big(\!\int_{0}^{T}u_{t}\,\#\,\mathrm{d}S_{t}\!\Big)G\!\right],$
for all $G\in\mathrm{Dom}(\delta)$. By density of $\mathrm{Dom}(\delta)$ in $L^{2}(\mathcal{A},\varphi)$, we identify $F-\varphi[F]=\delta(u)$, yielding (9). Uniqueness of $u$ follows from the free Itô isometry (51): if $\int_{0}^{T}(u_{t}-v_{t})\#\mathrm{d}S_{t}=0$ then $u=v$ in $L^{2}$. $\square$
@@ C.3  Free Time–Reversal of Semicircular Diffusions
The rigorous time–reversal step in Theorem 1 follows from Biane’s duality theorem for free stochastic integrals [37, Proposition 3.1]. For any adapted process $U_{t}\in L^{2}(\mathcal{A},\varphi)$ define
@@ D-A Voiculescu’s Free Entropy Functional
For a self-adjoint noncommutative random variable $X$ with spectral distribution $\mu$, Voiculescu’s free entropy $\chi(\mu)$ is defined by
@@ D-D Euler–Lagrange Equation and Free McKean–Vlasov Flow
The free entropy $\chi(\mu)$ can be regarded as the noncommutative analogue of the classical McKean–Vlasov free energy functional